# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ VI-A Feature attribution for MNIST image classification
In this subsection, the MNIST dataset, which is widely used for evaluating XAI methods [14, 29, 11], is considered to test the performance of EmSHAP. The MNIST dataset provides a standardized benchmark for assessing algorithm performance in tasks such as classification and pattern recognition, it consists of 70,000 grayscale images of digits (0-9), each having a size of 28x28 pixels, with 60,000 images allocated for training and 10,000 for testing.
@@
To classify the MNIST images, a simple CNN model is established, comprising two convolutional, two max-pooling layers and a fully connected output layer. The trained model achieved a classification accuracy of 97%, indicating a good performance. Given that CNN is a black-box model, it is difficult to understand the importance of different pixels. To perform this task, EmSHAP is employed to quantify the influence of input pixels on the classification results. It is trained and utilized to estimate the conditional probability distributions of the test data, which are further used to estimate the Shapley value.
@@
In order to enhance computational efficiency, we use average pooling to reduce the 28×28 image to 14×14 image, and the number of attributing features is reduced to 196. Fig 4 shows the feature attribution results using EmSHAP and 5 competitive methods.
Figure 4: Feature attributions on MNIST dataset
@@
As is shown in Fig 4, SVARM may face challenges in terms of interpretability when applied to image classification tasks, as it yields ambiguous results. Similarly, FastSHAP and VAEAC also produce ambiguous reconstructions. On the other hand, SamplingSHAP, KernelSHAP and EmSHAP all successfully attribute the digits of 3 and 8. A more delicate inspection, however, shows that SamplingSHAP and KernelSHAP yielded attribution that align with the classified digits, but tend to omit finer details. In contrast, EmSHAP exhibits finer details in feature attribution, especially for digit 3. The better performance is also verified by the higher SIC AUC-ADD value and lower SIC AUC-DEL value in Table III.
To test the performance of the proposed method under noisy data, Gaussian distributed noises with zero mean and varying degree of variances are added to the original images. The variances vary from 0.1 to 0.5, with an interval of 0.1. Feature attribution is then performed on these noisy images. Figs.5 and 6 present the attribution results produced by the six methods on the noisy images.
Figure 5: Feature attributions for noise-polluted image of digit 3. From top to bottom: noise variances of 0.1, 0.2, 0.3, 0.4, and 0.5.
Figure 6: Feature attributions for noise-polluted image of digit 8. From top to bottom: noise variances of 0.1, 0.2, 0.3, 0.4, and 0.5.
It can be seen from Figs.5 and 6 that when the variances of noise increase, the feature attribution performance of all methods deteriorates. Whilst the sharpest performance deterioration can be observed for SamplingSHAP, the performance for KernelSHAP and EmSHAP are relatively stable. This verifies the better robustness of EmSHAP in the case of noises.
VI-B Feature attribution for TREC question classification
The semantic classification of questions is a critical task in natural language processing, particularly in question-answering systems. Questions can be categorized into semantic classes based on their nature, such as inquiries about numbers, places, or objects. This classification aids in identifying potential answers. For instance, the question ”When did CNN begin broadcasting?” pertains to date and should be classified under the [Numeric] category.
In this subsection, the TREC question dataset, which includes six semantic classes, is used to evaluate the performance of EmSHAP. A CNN-based classifier (TextCNN) is used for classification [51], with a classification accuracy of 87.70%. For feature attribution, word-level attribution is performed to identify the trigger words that are most influential in determining the answer type.
Fig 7 presents a selection of questions randomly sampled from six distinct categories in the TREC test set. These questions are analyzed using the six methods to attribute word-level features to the trained TextCNN model. Compared to other methods, EmSHAP produces explanation results that better align with human language perception. The key advantages of EmSHAP over the other methods are summarized as follows: 1) Reduced attribution of generalized question words and question marks; 2) Increased attribution of key phrases; 3) Reduced attribution of connectives.
Figure 7: Feature attributions on TREC dataset
1) Reduced attribution of generalized question words and question marks: Certain key question words exhibit a strong correlation with question classification. For example, questions starting with ”Who” are classified as [Human] (question #4), those beginning with ”Where” as [Location] (question #5), and those starting with ”When” as [Numeric] (question #6). In these cases, the word-level feature attribution should reflect the prominence of these question words, meaning their Shapley values should be the highest. Conversely, the question word ”What” can initiate a wide range of question types (question #1–#3), making it weakly correlated with any specific category, and thus, less informative for question classification. Therefore, its Shapley value should be lower. Additionally, the presence of a question mark at the end of the sentence should not influence the classification, and as such, the Shapley value of the question mark should also be small.
As shown in Fig 7, for questions #1–#3, which begin with ”What,” EmSHAP reduces the contribution of ”What” and ”?” in feature attribution, focusing more on the contribution of the central phrase to question classification. This approach aligns with human cognitive processes. In contrast, other methods exhibit varying degrees of bias in attributing features to ”What” and ”?”.
2) Increased Attribution of Key Phrases: Compared to other methods, EmSHAP places greater emphasis on key phrases relevant to question classification. For instance, ”acronym NASA” corresponds to the category [Abbreviation], ”Phi Beta Kappa mean” corresponds to [Description], and ”popular toy in 1957” corresponds to [Entity].
3) Reduced Attribution of Connectives: Articles, prepositions, conjunctions, and similar words function to modify, connect, or supplement nouns or verbs, contributing to grammatical structures and the expression of complex meanings. Since these words do not stand alone in a sentence and are highly context-dependent, their contribution to question classification is relatively low in word-level feature attribution. Reducing the attribution of such words and prioritizing more essential short phrases is consistent with human semantic processing. As shown in Figure 7, EmSHAP reduces the attribution of words like ”the”, ”in” in [Entity] (question #3), ”in the” in [Human] (question #4), and ”are the” in [Location] (question #5).
VI-C Feature attribution for ADTI tabular classification
@@
Classifying adult income using census data is a typical tabular classification task and has been widely used to test XAI methods. In this subsection, we use this dataset to test the performance of EmSHAP on tabular data. The census data collects information on 48,842 independent individuals, involving 12 input variables related to individual income and a discrete output variable $y$, where $y=0$ means that the individual income is less than or equal to 50k, and $y=1$ means that the individual income is greater than 50k. After the individual data with missing values in the data are removed, 32,562 samples remain. 80% of these samples are randomly selected as the training set, and the remaining samples are used as the test set. To classify adult income, a three-layer MLP model is established, and the hidden layer size is set to 64. The model is trained using the training set and is optimized using the Adam optimizer with a learning rate of 0.001. The classification accuracy of the model can reach 84.75%, indicating that the model has good classification performance.
@@
To explain this predictive model, all six methods are used to quantify the contribution of input features. Fig 8 shows the mean absolute Shapley values of each feature.
Figure 8: Feature attributions for ADTI dataset
Among the six Shapley value estimation methods, SamplingSHAP, KernelSHAP, and EmSHAP identify age and capital gains as the most influential features in model predictions. This is in line with practical scenarios. Since the two variables are strongly correlated with income. For example, as people age, they typically accumulate more work experience and higher income. Capital gains are also positively correlated with income, as they generally stem from investment activities. In contrast, other factors like sex, marital status, and relationship tend not to be so related to income. On the other hand, SVARM also highlights relationship and sex as important. FastSHAP acknowledges that capital gains contribute most significantly, but the contributions of other variables are nearly identical. VAEAC suggests that marital status and relationship have a greater impact on adult income. Hence, in the case of ADTI dataset, the performance of feature attributions obtained from EmSHAP is also among the best.
@@
Figure 9: Feature attributions on ADTI dataset for different income classes
@@
To clearly illustrate the contribution of each feature to model classifications, we have categorized the data into two groups: income $\leq 50$K and income $>50$K. Fig. 9 presents the mean absolute Shapley values for each feature across the two income categories. In Fig. 9, bar charts with slashes and circles represent the absolute average Shapley values for the classification of individuals with income $\leq 50$K and $>50$K, respectively. It can be observed that, for the income $\leq 50$K group, age has the greatest contribution to the model’s classification, while for the income $>50$K group, capital gains hold the greatest influence. This difference can be explained. As individuals with income $\leq 50$K are typically younger or have less work experience, by accumulating age and work experience, their income will increase. Therefore, in the low-income group, age plays a more significant role in feature attribution. On the other hand, individuals with higher incomes tend to have stronger investment capabilities and more available funds for investment, which results in higher capital gains. Consequently, capital gains are a stronger attributor for individuals with income $>50$K.
VI-D Feature attributions for ETT dataset
@@
In this subsection, the electricity transformer dataset (ETT), which has been widely used in long-term and short-term time series prediction tasks [52, 53], is considered. The time series data in ETT dataset is typically non-stationary and contains a reasonable portion of outliers. The dataset consists of 6 power load features and an output feature of ”oil temperature”, sampled at different temporal granularities. Here only the hourly data ETTh1 is used, which contains a total of 17420 samples.
@@
In the ETT dataset, the task is to make predictions of oil temperature based on the 6 power load features. This will help the operators to make corrective operations before the oil temperature gets out of the normal interval. For the purpose of constructing predictive model, the Informer model [53] is applied, with the first 80% of the dataset used as training data and the last 20% as test data. The Informer model performs well in the test data, achieving a mean square error of less than 0.02. Despite the high accuracy, it is still desirable to know which input feature is more response for the predictions, which can be determined through estimation of Shapley values using EmSHAP and the other 5 competitive methods.
@@
The feature attribution results can be referred to in Table III, which shows that EmSHAP outperforms competitive methods in terms of the SIC AUC metric. For a clear display, Fig. 10 presents the heatmap of estimated Shapley values using EmSHAP for the test set. It can be seen from Fig. 10 that the first four variables, HUFL, HULL, MULL, and MULL are more responsible for the model predictions, whilst the contributions of LUFL and LULL are negligible. This is in accordance with the feature attribution results of other competitive methods.
@@
Figure 10: Heatmap of estimating Shapley values using EmSHAP in the ETTh1 prediction task
@@
In order to further test the robustness of EmSHAP under adversarial attacks, a series of adversarial samples with varying attack amplitudes were generated using the Fast Gradient Sign Method (FGSM) [54] and added to the original test data. Based on the new test data, feature attribution is again performed using EmSHAP, FastSHAP and VAEAC, and the results are shown in Table IV. The reason for not comparing with other methods is that they do not involve a training stage. It can be seen from Table IV that the SIC AUC-ADD values of all three methods decrease after adding the adversarial samples. A similar result can be observed for the SIC AUC-DEL values, which increased for all the three methods, indicating the deterioration of feature attribution performance. A more delicate observation on the SIC AUC-ADD and SIC AUC-DEL values, however, shows that EmSHAP consistently outperforms the other two methods across all tested amplitudes.
TABLE IV: SIC AUC results for different attack amplitude on ETT dataset
Explainer SIC AUC Attack amplitude

### v4

@@ VI-A Feature attribution for MNIST image classification
In this subsection, the MNIST dataset, which is widely used for evaluating XAI methods [12, 36, 13], is considered to test the performance of EmSHAP. The MNIST dataset provides a standardized benchmark for assessing algorithm performance in tasks such as classification and pattern recognition. It consists of 70,000 grayscale images of digits (0-9), each having a size of 28x28 pixels, with 60,000 images allocated for training and 10,000 for testing.
@@
To classify the MNIST images, a simple CNN model is established, comprising two convolutional, two max-pooling layers, and a fully connected output layer. The trained model achieved a classification accuracy of 97%, indicating a good performance. Given that CNN is a black-box model, it is difficult to understand the importance of different pixels. To perform this task, EmSHAP is employed to quantify the influence of input pixels on the classification results. It is trained and utilized to estimate the conditional probability distributions of the test data, which are further used to estimate the Shapley value.
@@
In order to enhance computational efficiency, we use average pooling to reduce the 28×28 image to a 14×14 image, and the number of attributing features is reduced to 196. Fig 5 shows the feature attribution results using EmSHAP and seven competitive methods.
Figure 5: Feature attributions on MNIST dataset.
@@
As is shown in Fig 5, SVARM and SHAP-IQ may face challenges in terms of interpretability when applied to image classification tasks, as they yield ambiguous results. Similarly, FastSHAP and VAEAC also produce ambiguous reconstructions. On the other hand, SamplingSHAP, KernelSHAP, ST-SHAP, and EmSHAP all successfully attribute the digits 3 and 8. A more delicate inspection, however, shows that SamplingSHAP, KernelSHAP, and ST-SHAP yielded attribution that align with the classified digits, but tends to omit finer details. In contrast, EmSHAP exhibits finer details in feature attribution, especially for digit 3. The better performance is also verified by the higher SIC AUC-ADD value and lower SIC AUC-DEL value in Table II. The robustness of different methods is also tested by adding Gaussian noise with standard deviations varying from 0.1 to 0.5, with an interval of 0.1. The results are shown in Appendix D, which verifies the better robustness of EmSHAP in the case of noises.
VI-B Feature attribution for ADTI tabular classification
@@
Classifying adult income using census data is a typical tabular classification task and has been widely used to test XAI methods. In this subsection, we use this dataset to test the performance of EmSHAP on tabular data. The census data collects information on 48,842 independent individuals, involving 12 input variables related to individual income and a discrete output variable $y$, where $y=0$ means that the individual income is less than or equal to 50k, and $y=1$ means that the individual income is greater than 50k. After the individual data with missing values are removed, 32,562 samples remain. 80% of these samples are randomly selected as the training set, and the remaining samples are used as the test set. To classify adult income, a three-layer MLP model is established, and the hidden layer size is set to 64. The model is trained using the training set and is optimized using the Adam optimizer with a learning rate of 0.001. The classification accuracy of the model can reach 84.75%, indicating that the model has good classification performance.
@@
To explain this predictive model, all eight methods are used to quantify the contribution of input features. Fig 6 shows the mean absolute Shapley values of each feature.
Figure 6: Feature attributions for ADTI dataset.
Among the eight Shapley value estimation methods, SamplingSHAP, SHAP-IQ, KernelSHAP, and EmSHAP consistently identify age and capital gains as the most influential features in income prediction, which aligns with real-world intuition. These two variables exhibit strong correlations with income: older individuals typically accumulate more work experience and thus higher earnings, while capital gains are directly associated with investment activities. In contrast, demographic attributes such as sex, marital status, and relationship generally show weaker correlations with income. SVARM diverges from this trend by assigning higher importance to relationship and sex, whereas ST-SHAP fails to effectively distinguish the most critical input features. FastSHAP correctly emphasizes the dominant role of capital gains, but treats the remaining variables as having nearly uniform contributions. Similarly, VAEAC attributes substantial importance to marital status and relationship, suggesting a different explanatory perspective. Overall, for the ADTI dataset, EmSHAP produces feature attributions that are not only consistent with domain knowledge but also rank among the most reliable explanations across the evaluated methods.
@@
Figure 7: Feature attributions on ADTI dataset for different income classes.
@@
To clearly illustrate the contribution of each feature to model classifications, we have categorized the data into two groups: income $\leq 50$K and income >50>50K. Fig. 7 presents the mean absolute Shapley values for each feature across the two income categories. In Fig. 7, bar charts with slashes and circles represent the absolute average Shapley values for the classification of individuals with income $\leq 50$K and >50>50K, respectively. It can be observed that, for the income $\leq 50$K group, age has the greatest contribution to the model’s classification, while for the income >50>50K group, capital gains hold the greatest influence. This difference can be explained. As individuals with income $\leq 50$K are typically younger or have less work experience, by accumulating age and work experience, their income will increase. Therefore, in the low-income group, age plays a more significant role in feature attribution. On the other hand, individuals with higher incomes tend to have stronger investment capabilities and more available funds for investment, which results in higher capital gains. Consequently, capital gains are a stronger attribution for individuals with income >50>50K.
VI-C Feature attributions for ETT dataset
@@
In this subsection, the electricity transformer dataset (ETT), which has been widely used in long-term and short-term time series prediction tasks [65, 66], is considered. The time series data in the ETT dataset is typically non-stationary and contains a reasonable portion of outliers. The dataset consists of 6 power load features and an output feature of ”oil temperature”, sampled at different temporal granularities. Here, only the hourly data ETTh1 is used, which contains a total of 17420 samples.
@@
In the ETT dataset, the task is to make predictions of oil temperature based on the 6 power load features. This will help the operators to make corrective operations before the oil temperature gets out of the normal interval. For the purpose of constructing a predictive model, the Informer model [66] is applied, with the first 80% of the dataset used as training data and the last 20% as test data. The Informer model performs well in the test data, achieving a mean square error of less than 0.02. Despite the high accuracy, it is still desirable to know which input feature is more response for the predictions, which can be determined through estimation of Shapley values using EmSHAP and the other seven competitive methods.
@@
The feature attribution results can be referred to in Table II, which shows that EmSHAP outperforms competitive methods in terms of the SIC AUC metric. For a clear display, Fig. 8 presents the heatmap of estimated Shapley values using EmSHAP for the test set. It can be seen from Fig. 8 that the first four variables, HUFL, HULL, MULL, and MULL, are more responsible for the model predictions, whilst the contributions of LUFL and LULL are negligible. This is in accordance with the feature attribution results of other competitive methods.
@@
Figure 8: Heatmap of estimating Shapley values using EmSHAP in the ETTh1 prediction task.
@@
In order to further test the robustness of EmSHAP under adversarial attacks, a series of adversarial samples with varying attack amplitudes is generated using the Fast Gradient Sign Method (FGSM) [67] and added to the original test data. Based on the new test data, feature attribution is again performed using EmSHAP, FastSHAP, and VAEAC, and the results are shown in Table III. The reason for not comparing with other methods is that they do not involve a training stage. It can be seen from Table III that the SIC AUC-ADD values of all three methods decrease after adding the adversarial samples. A similar result can be observed for the SIC AUC-DEL values, which increased for all three methods, indicating the deterioration of feature attribution performance. A more delicate observation on the SIC AUC-ADD and SIC AUC-DEL values, however, shows that EmSHAP consistently outperforms the other two methods across all tested amplitudes.
TABLE III: SIC AUC results for different attack amplitudes on the ETT dataset
Explainer
SIC
AUC
Attack amplitude
@@
FastSHAP
[…]
VAEAC
[…]
EmSHAP
[…]
VI-D Feature attributions for MIMIC-III
The MIMIC-III dataset [68] is a large-scale, real-world benchmark in medical machine learning, containing clinical documents with International Classification of Diseases (ICD-9) annotations from 2001 to 2012. In this subsection, we focus on the automated medical coding task [69] by using 11,371 discharge summaries(each truncated to 2,500 words) to predict the 50 most frequent ICD-9 codes in a multi-label setting. The data preprocessing procedure is followed by [70], and the data are split into 8,067/1,574/1,730 samples for training, validation, and testing, respectively. We adopt the Convolutional Attention for Multi-Label (CAML) classification model [70], which achieves a top-5 precision of 0.70 on this large-scale task. To interpret its black-box predictions, EmSHAP is applied for word-level attribution and compare it with seven attribution baselines, as reported in Table II.
Figure 9: Feature attributions on MIMIC-III dataset.
Due to space limitations, Fig. 9 presents a partial excerpt of the medical record with ID 120589 together with its feature attribution results, more results are reported in Appendix D. In Fig. 9, the top two medical codes predicted by CAML are ’584.9’ (acute renal failure, unspecified) and ’428.0’ (congestive heart failure, unspecified), which are consistent with the real codes. In the visualization, red and green highlight positive and negative Shapley attributions, respectively, with deeper colors indicating stronger contributions. The results show that EmSHAP successfully identifies clinically meaningful terms such as “renal”, “failure”, “endocarditis”, “pleuritic”, and “chest”, which align with medical knowledge and confirm that the model predictions rely on relevant clinical cues. At the same time, EmSHAP is not without limitations: due to the complexity and high dimensionality of medical text, irrelevant connective words such as “and”, “from”, and “to” are occasionally assigned spuriously high attributions. This phenomenon may be caused by the GRU’s sensitivity to frequent tokens and the amplification of contextual dependencies during conditional probability modeling. Nevertheless, as indicated by the explanation metric SIC AUC in Table II, EmSHAP consistently highlights key clinical features, thereby providing more interpretable and trustworthy results than existing methods in the automated medical coding task.
VI-E Performance analysis of EmSHAP under different model structures and parameter settings
In this subsection, the performance of EmSHAP under different model structures is compared and analyzed.
VI-E1 Dynamic masking vs fixed-rate masking
As an important component of EmSHAP, the dynamic masking mechanism enabling the model’s performance. To better show the functionality of dynamic masking, the performance of EmSHAP under dynamic masking and fix-rate masking is compared, and the results are shown in Appendix E (Table E1). As shown in Table E1, the dynamic masking achieves better performance than fixed-rate masking for all cases. In addition, using the fixed-rate masking requires manually tuning the optimal masking rate, which is undesired in practice. In contrast, the dynamic masking mechanism allows the model to learn masking rate automatically, which leads to better performance [71].
VI-E2 Analysis of context vector
In addition to investigating the role of the context vector $\bm{\gamma}$ in EmSHAP, we analyze how the dimension of $\bm{\gamma}$ affects the model performance. The results of these experiments are presented in Appendix E (Table E2). As shown in Table E2, increasing the context vector dimension leads to a steady improvement in the SIC AUC metrics across all datasets. However, since a larger dimension also increases model complexity, choosing an appropriate dimension is essential to balance efficiency and accuracy.
VI-E3 Trade-off between accuracy and efficiency
The trade-off between accuracy and efficiency is a key consideration in model design. In EmSHAP, increasing the latent dimension enhances predictive performance but also raises computational cost in terms of training and inference time, memory usage, and energy consumption. The performance of EmSHAP under different latent dimensions on four datasets is presented in Appendix E (Table E3). As shown in Appendix E (Table E3), enlarging the latent dimension from 16 to 128 yields consistent gains in SIC AUC, accompanied by higher resource demands. These findings underscore the need to balance accuracy with computational efficiency when selecting latent dimensions for practical deployment.
VI-E4 Comparison of different RNN-based proposal networks
To further investigate the impact of different RNN-based architectures, we compared GRU with LSTM, BiGRU, and attention-based GRU (Att-GRU). The results, summarized in Appendix E (Table E4), indicate that although LSTM, BiGRU, and Att-GRU occasionally achieve slightly higher SIC AUC scores (e.g., on MNIST and ADTI), these gains come at the expense of substantially higher runtime and memory consumption. Given that the primary goal of EmSHAP is to provide accurate yet efficient explanations, GRU offers the most favorable trade-off between performance and computational cost. Moreover, EmSHAP remains model-agnostic: the GRU module can be readily replaced with LSTM, BiGRU, Att-GRU, or other advanced sequence encoders if additional computational resources are available.
VI-E5 Ablation study
An ablation study is performed to reflect the performance of EmSHAP on different standardized models, which involves the combinations of vanilla EBM, the GRU network, and the dynamic masking mechanism. The detailed results are reported in Appendix E (Table E5), which presents both feature attribution performance and runtime across different model configurations. As shown in Table E5, although EmSHAP requires more time for training, it demonstrates superior performance in terms of the SIC AUC metric. More specifically, the vanilla EBM yields the lowest performance, but by gradually adding the dynamic masking mechanism and GRU network, the performance increases. This clearly shows that the combined effect of these mechanisms contributes to the better performance of EmSHAP.
VII Conclusion
This paper proposes an energy-based model for Shapley value estimation (EmSHAP) for feature attribution in deep learning-based model. In EmSHAP, an energy-based model (EBM) is used to estimate the Shapley contribution function under different feature combinations. The EBM consists of a ResNet structure for approximating the energy function and a GRU network for estimating the partition distribution. By mapping the input features into a latent space using the GRU network, the influence of feature ordering on the calculation of the contribution function can be eliminated. In order to explore a wider range of feature combinations, a dynamic masking mechanism is introduced during training, which empowers EmSHAP with better estimation accuracy and generalization capability. In order to evaluate the performance of the proposed method, detailed theoretical analysis on error bound and extensive application studies are conducted, which show that EmSHAP has a tighter error bound and exhibits higher accuracy compared with state-of-the-art approaches.
Although EmSHAP provides a theoretically grounded and empirically competitive approach to Shapley-based feature attribution, several limitations still remain: i) the GRU architecture may be suboptimal for certain data modalities such as graph-structured or highly sparse features, highlighting the need for more flexible conditional modelling; ii) the quality of the proposal distribution directly affects conditional density estimation, the accuracy and stability may not be that good for hard-to-model feature subsets; iii) EmSHAP has a computational complexity that is proportional to the quadratic of input dimensions, hence it may need additional computational resources to handle high dimensional data sets; iv) the performance of EmSHAP may deteriorate under non-stationary data. Future research direction can be focused on: i) incorporating causal attribution methods to disentangle spurious correlations and generate counterfactual explanations; ii) designing adaptive sampling strategies that dynamically prioritize these hard-to-model subsets to improve the quality of the proposal distribution; iii) dimensionality-aware extensions to enhance applicability in high-dimensional domains; and iv) development of time-varying modelling approaches to handle non-stationary data.
Appendices
VII-A Error Bound of KernelSHAP
KernelSHAP believes that the calculation of the Shapley value can be regarded as an additive model, that is, the contribution function $v(S)$ of the variable set $S$ can be approximated by the sum of the variable weights within $S$,
$\small v(S)\approx\beta_{0}+\sum_{i\in S}\beta{i},$ (A1)
where $\beta_{0}$ is the mean of $v(S)$, $\beta_{i}$ is the weight scalar of the $i$-th variable. By designing a linear regression model and solving it using the weighted least square method, the weight of each variable can be obtained. The optimization formula of KernelSHAP is as follows.
$\small\mathop{\min}_{\beta_{0},...,\beta_{|D|}}\sum_{S\in D}\psi(S)\left(\beta_{0}+\sum_{i\in S}\beta_{i}-v(S)\right)^{2},$ (A2)
where $\bm{\beta}=[\beta_{0},...,\beta_{d}]^{T}\in\mathbb{R}^{|D|+1}$ is the weight of each variable, $\psi(S)\in[0,1]$ is called the Shapley kernel. To solve the above optimization problem, Covert et al. [14] introduced an binary vector $\mathbf{b}\in\mathbb{R}^{|D|+1}$, $b_{i}=1$ indicates the $i$-th variable is in the subset $S$, and otherwise $b_{i}=0$. Thus, Eq.(A2) can be rewritten as follows.
The matrix form of Eq.(A3) is as follows.
where $K$ is the number of sampling times, $\mathbf{B}\in\mathbb{R}^{(|D|+1)\times K}$ is a binary matrix indicating which variables are in $S$, $\bm{\psi}\in\mathbb{R}^{K\times K}$ is a diagonal matrix with Shapley kernel in its diagonal, $\mathbf{v}\in\mathbb{R}^{K\times 1}$ is the contribution function vector of different samples.
Eq.(A4) can be solved by weighted least square method to obtain the estimation $\hat{\bm{\beta}}$, by taking the first-order derivative of $\bm{\beta}$ in Eq.(A4) and setting it to zero, the derivation of weighted least square is given by,
For the details of KernelSHAP, interested readers can refer to Refs. [13, 14].
Let $\hat{\bm{\Sigma}}=\frac{1}{K}\mathbf{B}\bm{\psi}\mathbf{B}^{T}$ and $\hat{\bm{\Gamma}}=\frac{1}{K}\mathbf{B}^{T}\bm{\psi}\mathbf{v}$, $\mathbb{E}[\hat{\bm{\Sigma}}]=\bm{\Sigma}^{*}$ and $\mathbb{E}[\hat{\bm{\Gamma}}]=\bm{\Gamma}^{*}$. Assuming that $\hat{\bm{\Sigma}}$ and $\bm{\Sigma}^{*}$ is invertible, we have $\hat{\bm{\beta}}=\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}$, $\bm{\beta}^{*}=\mathbb{E}[\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}$]. The error bound of KernelSHAP can be described as the mean absolute deviation between the estimated contribution function $\hat{v}(S)$ and the true contribution function $v(S)$, this error bound is also related to the $\hat{\bm{\beta}}$ and $\bm{\beta}^{*}$ as follows.
[…]
The decomposition of Eq.(A6) is given by,
$\qquad\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ (A7)
[…]
where $||\cdot||$ is the $L_{2}$ norm and $\text{Cov}(\cdot,\cdot)$ is the covariance. Next we introduce three corollaries and show that $\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\leq\epsilon_{3}^{2}$ in some specific cases, where $\epsilon_{3}$ is a small positive number.
Corollary A1.
(Concentration of $\hat{\bm{\Sigma}}$) For all $\epsilon_{3}>0$, we have,
[…]
Proof. Noting that the elements in $\mathbf{B}$ are 0 or 1 and the elements in $\bm{\psi}$ is bounded with $[0,1]$. Thus the element in $\hat{\bm{\Sigma}}$ and $\bm{\Sigma}^{*}$ also bounded with $[0,1]$. According to Hoeffdings inequality, each element in $\hat{\bm{\Sigma}}$ is bounded with their expectation values, that is, for all $\epsilon_{3}>0$
where, $\mathbb{E}[\hat{\bm{\Sigma}}_{ii}]=\frac{1}{2}$ and $\mathbb{E}[\hat{\bm{\Sigma}}_{ij}]=\alpha$, so
$\small\begin{cases}\mathbb{P}(|\frac{1}{K}\sum_{k=1}^{K}(\hat{\bm{\Sigma}}_{ii}-\frac{1}{2})|\geq\epsilon_{3})\leq 2\exp(-2K\epsilon_{3}^{2})\\ \mathbb{P}(|\frac{1}{K}\sum_{k=1}^{K}(\hat{\bm{\Sigma}}_{ij}-\alpha)|\geq\epsilon_{3})\leq 2\exp(-2K\epsilon_{3}^{2}).\end{cases}$ (A10)
By using Booles inequality [72], Eq.(A8) can be obtained.
Corollary A2.
(Concentration of $\hat{\bm{\Gamma}}$)For all $\epsilon_{3}>0$, we have,
[…]
Proof. According to Assumption 4, the contribution value of any subset $S$ is bounded within $[0,1]$, that is $v(S)\in[0,1]$. Then $\mathbf{v}$ is also bounded within $[0,1]$, which indicates $\bm{\Gamma}^{*}\in[0,1]$, by using Hoeffdings inequality, each element in $\hat{\bm{\Gamma}}$ is bounded with their expectation values, that is, for all $\epsilon>0$
By using Booles inequality [72], the concentration of $\hat{\bm{\Gamma}}$ can be obtained.
Corollary A3.
(Concentration of $\hat{\bm{\Sigma}}^{-1}$) For all $\epsilon_{3}>0$, we have,
[…]
Corollary A4.
For the fixed input dimension $|D|$, for any small positive number $\epsilon_{3}$, there exists a large $K$ that